Circular No. 8047

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF NEPTUNE M. Holman, Harvard-Smithsonian Center for Astrophysics (CfA); J.

Kavelaars, National Research Council of Canada; T. Grav, University of Oslo and CfA; and W. Fraser and D. Milisavljevic, McMaster University, report the discovery of three satellites of Neptune on CCD images obtained in Aug. 2002 with the 4-m Blanco telescope at Cerro Tololo. Team members B. Gladman, J.-M. Petit, P. Rousselot, and O. Mousis conducted followup observations with the European Southern Observatory 8.2-m UT3 telescope, as did Grav with the 2.6-m Nordic Optical Telescope and P. Nicholson, Gladman, and V. Carruba with the 5-m Hale Telescope at Palomar; B. G. Marsden and R. Jacobson provided orbital predictions. The astrometry, with orbital elements (assuming identity with single-night observations made in 2001 with the 3.6-m Canada-France-Hawaii Telescope at Mauna Kea and the Blanco telescope) and ephemerides by Marsden, appear on MPEC 2003-A75: S/2002 N 1, a = 0.147 AU, e = 0.43, i = 121^{o} (to the ecliptic), H = 9.7; S/2002 N 2, a = 0.135, e = 0.17 (assumed), i = 57^{o}, H = 11.0; S/2002 N 3, a = 0.143, e = 0.47 (assumed), i = 43^{o}, H = 10.6.

NOVA IN NGC 185 M. Papenkova, J. Graham, and W. Li, University of California at

Berkeley, report the LOTOSS/KAIT (cf. IAUC 7906) discovery of an apparent nova located at α = 0^{h}39^{m}06^{s}.07, δ = +48^{o}22′56′′.4 (equinox 2000.0), which is 81′′.4 east, 162′′.0 north of the center of NGC 185. Available photometry: 2002 Dec. 2 UT, [19.0; 11, 19:; 2003 Jan. 5.1, 18.4; 6.1, 18.5.

SUPERNOVA 2003H IN NGC 2207 M. Hamuy, Carnegie Observatories, reports on spectroscopic observa-

tions made (as detailed on IAUC 8045) by J. Maza on Jan. 9 and 10 UT: “The spectrum of SN 2003H (cf. IAUC 8045, 8046) shows a red continuum with a prominent absorption at 575.1 nm; if this is He I 587.6-nm, the expansion velocity is 9100 km/s. A more shallow blend of absorption lines can be seen at 600–670 nm, and an isolated absorption is centered at 690 nm. The Ca II triplet displays a prominent P-Cyg profile. This spectrum has some similarity to that of the type-Ib supernova 1984L before maximum light, but the continuum is much more red. There is no evidence for strong Na I D interstellar lines, so dust extinction must be small and the red color is probably intrinsic to SN 2003H.”

2003 January 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green